OpenStudy (anonymous):

What is the solution of sqrt 5x+1 - sqrt x = 5 ?

OpenStudy (anonymous):

\[\sqrt{5x+1} - \sqrt{x} = 5\]

OpenStudy (kewlgeek555):

So sorry for the late reply. I have more than two people waiting.

OpenStudy (anonymous):

It's okay I figured out the answer ( :

OpenStudy (kewlgeek555):

Oh. Sorry, a medal to you then. ;]

OpenStudy (anonymous):

Thanks ( :